# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v1

@@ Abstract
Introduction — Soon after the discovery of high temperature superconductivity (HTS) in a family of bilayer nickelates – first in pressurized bulk crystals [1, 2, 3, 4], and later in compressively strained films [5, 6, 7, 8] – intense effort has been devoted to understanding the emergent properties of these systems. Experiments have begun to shed some light on structural properties [9, 10], electronic band structure [11, 12, 13, 14] and magnetic orders proximate to superconductivity [15, 16, 17, 18, 19, 20, 21]. On the theoretical front, an impressive collection of works have already provided detailed proposals for the electronic structure, correlations and their implications for superconductivity [22, 23, 24, 25, 26, 11, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]
In the wake of a new experimental discovery, there is the temptation towards classification, i.e. to ascertain whether the new phenomena belongs to a category of its own, or represents an interesting variation on a known theme. While HTS in the thin film bilayer nickelate is reminiscent of the overdoped cuprates [55], there remain obvious distinctions, the most prominent one being the discrepancy in valence count. In, for example, the La-based cuprate parent compound La2CuO4, the low energy behavior is determined by the copper $d^{9}$ configuration, in which electronic excitations stem from the $d_{x^{2}-y^{2}}$ orbital and oxygen $p$ orbitals [56, 57, 58]. By contrast, the $d^{7.5}$ valence in the bilayer nickelate La3Ni2O7 necessitates the importance of multiple $d$-orbitals and Hund’s coupling. To determine whether such differences in local quantum chemistry lead to new emergent behavior, it is useful to take a step back and ask questions that are more qualitative in nature.
@@
Bilayer two-orbital Hubbard model — The bilayer nickelate consists of a $d^{7.5}$ configuration of Nickel electrons in a tetrahedral crystalline environment. Consequently, the low energy electronic excitations come from partially filled, non-degenerate Nickel $d_{3z^{2}-r^{2}}$ (denoted as ``$Z"$) and $d_{x^{2}-y^{2}}$ (denoted as ``$X"$) orbitals. Thus, we begin with a two-orbital Hubbard model on a bilayer square lattice (see Fig. 1(a)), incorporating the $X,Z$ orbitals of Ni ions. Owing to the shape and orientation of these orbitals, it follows that electrons in the $Z$-orbital tunnel predominantly between the Nickel planes of the bilayer, whereas electrons in the $X$-orbital primarily tunnel within these planes (see Fig 1(a)). The Hamiltonian that incorporates only these strongest orbital hybridization effects is
[…]
@@
Magnetism in the insulating $d^{8}$ limit — At $x=0$, both the $X$ and $Z$-orbitals are half-filled and the model in Eq.( 1) reduces to an effective two orbital, spin-1/2 bilayer Heisenberg model (see Fig. 2):
@@
Figure 3: (a) Phase diagram from bond-operator mean-field theory (BOMFT). Blue, yellow region denotes interlayer singlet and Néel order, respectively. The black line indicates the second order phase transition boundary between these two phases. (b) $J_{H}$ dependence of orbital-summed Néel order, $N=N_{x^{2}-y^{2}}+N_{z^{2}}$ (purple line) from Holstein-Primakov theory at $J_{\perp}=5$. Orbital-resolved Néel orders, $N_{x^{2}-y^{2}}=(-1)^{i+l}\langle S^{d;z}_{i,l}\rangle$ and $N_{z^{2}}=(-1)^{i+l}\langle S^{f;z}_{i,l}\rangle$, are plotted in blue, red dashed line, respectively. The system transitions from low-spin (spin-1/2) AFM on the $X$-orbital at small $J_{H}$ to high-spin (spin-1) AFM from both orbitals at large $J_{H}$.
@@
Within the BOMFT, one begins by assuming an interlayer singlet ground state and describes the system using a local singlet–triplet basis, which corresponds to the ground and first excited states in the $J_{\parallel}\rightarrow 0$ limit. The critical interlayer coupling $J_{\perp}^{c}$ is then determined by the point at which the triplet excitation gap closes, due to the finite in-plane coupling $J_{\parallel}$ effects. Using two variational parameters—the condensation of interlayer singlet $\langle s\rangle$ and the chemical potential $\mu$—we numerically solve the self-consistent Hamiltonian and determine the phase boundary and its analytic form [65]. Figure 3(a) shows the resulting phase diagram from BOMFT. As expected, $J_{\perp}^{c}$ decreases with increasing $J_{H}$ and approaches $J_{\perp}^{c}\rightarrow 6.098$ in the $J_{H}\rightarrow\infty$ limit, in closer agreement with the QMC result $J_{\perp}^{c}=7.15$ for the bilayer spin-1 Heisenberg model [61].
@@
Next, to understand the spin-wave spectrum deep inside the antiferromagnetic phases, we use the standard Holstein–Primakov (HP) spin-wave theory [62, 66]. We evaluate the orbital-resolved Néel order $N_{x^{2}-y^{2}}=(-1)^{i+l}\langle S^{d;z}_{i,l}\rangle$ and $N_{z^{2}}=(-1)^{i+l}\langle S^{f;z}_{i,l}\rangle$ and their sum, $N\equiv N_{x^{2}-y^{2}}+N_{z^{2}}$. In Fig. 3(b), we plot $J_{H}$ dependence of $N$, $N_{x^{2}-y^{2}}$ and $N_{z^{2}}$ at a representative value $J_{\perp}=5$, where the AFM phase is expected in all range of $J_{H}$. For small $J_{H}$, $N_{z^{2}}$ remains suppressed and only $N_{x^{2}-y^{2}}$ exhibits finite magnetic order. This corresponds to a low-spin (spin-1/2) antiferromagnetic state. As $J_{H}$ increases, Hund’s coupling more transmits $J_{\perp}$ to the $X$-orbital, suppressing $N_{x^{2}-y^{2}}$, while simultaneously enhancing the transmission of $J_{\parallel}$ to the $Z$-orbital, thereby increasing $N_{z^{2}}$ and $N$. In the large-$J_{H}$ limit, the two order parameters converge to the same value, indicating a high-spin (spin-1) antiferromagnetic state. Even in the $J_{H}\to\infty$ limit with a small $J_{\perp}$, $N\to 0.80$, matching the $S=1$ HP result. In SM, we present a full map of the order parameters across the $(J_{\perp},J_{H})$ space [65]. Thus, we have used several complementary approaches to establish that magnetism in the $d^{8}$ insulating regime can take on either high-spin, low spin forms, or can exhibit a spin-gapped, non magnetic ground state.
@@
Figure 4: Hartree-Fock (HF) results for Néel order ($N$) under hole doping $x$ of (a) low-spin and (b) high-spin configurations for values of $U/t_{\parallel}$. Here, $N=(-1)^{i+\ell}\langle S^{d;z}_{i,\ell}+S^{f;z}_{i,\ell}\rangle$ is an orbital-summed Néel order. We used the hopping parameters $t_{\parallel}=0.483$, $t_{\perp}=0.635$, and onsite energy difference $\Delta=0.367$ from DFT calculations [22]. In both cases, we observe first-order or weakly first-order transitions. The AFM order in the high-spin case is more robust than in the low-spin case, as indicated by a larger critical doping $x_{c}$
We present the HF results for an intermediate $U=4t_{\|}$ in Fig.4. In the low spin case (Fig.4(a)), when the two orbitals are completely decoupled ($J_{H}=0$), the HF result coincides with the single layer single orbital square lattice Hubbard model, where for $U=4t$, AFM order vanishes at $x_{c}\approx 0.207$ through a first order phase transition [67]. Such first order antiferromagnetic phase transitions at $T=0$ are typical within the HF approximation. With a larger $J_{H}=0.4U$, the low spin AFM order gets suppressed compared to the decoupled limit: the spin order survives to a smaller $x_{c}\approx 0.18$ and its magnitude also becomes smaller. By contrast, the high spin AFM can survive to a much larger hole doping, as can be seen in Fig.4(b). Moreover, the nature of the high spin AFM phase transition as a function of $x$ is also different from the low spin counterpart. We found that at large enough $J_{H}$, the high spin AFM order parameter $N$ can exhibit a discontinuous jump to a smaller value, but after this it appears to continuously decrease as $x$ increases further. This behavior seems to be a combination between both a first order and a continuous phase transition.
@@
Discussion — We have analyzed parent compounds of bilayer nickelate with $d^{8}$ electron configuration and have shown that there are several distinct types of magnetic behavior depending on the interplay between Hund’s coupling ($J_{H}$) and superexchange ($J_{\bot},J_{\|}$). We have identified a crossover from a low-spin AFM order at small $J_{H}$ to a high-spin AFM order at larger $J_{H}$. We have also identified a spin-gapped insulating state that occurs when $J_{\perp},J_{H}\gg J_{\parallel}$. The nature of magnetism in the parent compound will certainly have implications for superconductivity.
We have shown here that the low energy physics of the doped low-spin state is adequately described by a single orbital bilayer Hubbard model when $J_{H}/J_{\perp}$ is vanishingly small. When $J_{H}/J_{\perp}$ is small, but finite, an exchange coupling between the two layers will be generated by virtual processes. In contrast, the low energy effective description of the doped high spin state is a bilayer-Kondo lattice with ferromagnetic Kondo coupling due to Hund’s coupling and a weak interlayer exchange coupling. If one considers the extreme limit where $J_{H}\rightarrow\infty$, the Kondo model can be further simplified to a bilayer type-II $t$$J$ model, as two of us derived in the previous paper [24]. We stress that all three different regimes can be reconciled by a double-Kondo model  [60].
Within mean-field theory, we found that upon hole doping, the high-spin AFM is more robust. It remains unclear at present whether such robustness suggests that the high-spin state more strongly competes with superconductivity, or provides a stronger pairing ‘glue’ than the low-spin counterpart. Lastly, the spin gapped insulator, when doped exhibits superconducivity rather distinct than either the high- or low-spin states, with strong interalyer pairing as was shown recently [59, 60].
@@
Our findings are pertinent to recent experimental studies of superconductivity in compressively strained thin films of the bilayer. With compressive strain, we naturally expect that the in-plane superexchange will be enhanced whereas with tensile strain, the opposite would hold. Given the rich magnetic phase diagram revealed in our study, the experimental observation of distinct spin states in strain-tuned systems represents a promising arena for future exploration.
@@ 5 Departments of Applied Physics, Stanford University, Stanford, CA 94305, USA
[…]
@@
[…]
@@
In the $J_{\parallel}\rightarrow 0$ limit, we can define local singlet and triplet states. In our model, each site contains four $S=1/2$ moments—two orbitals on each of the two layers—which decompose into irreducible spin multiplets as $\bm{1/2}\otimes\bm{1/2}\otimes\bm{1/2}\otimes\bm{1/2}=2\times\bm{0}\oplus 3\times\bm{1}\oplus 1\times\bm{2}$. That is, the local Hilbert space consists of two singlets, nine triplet states, and five quintet states. While the detailed structure of these multiplets depends on the ratio $r=J_{H}/J_{\perp}$, for all values of $r$, the lowest-energy state is a singlet, denoted by |$\ket{s}\equiv s^{\dagger}\ket{0}$, and the first excited states are the three triplet modes |$\ket{t_{m}}\equiv t^{\dagger}_{m}\ket{0}$ with $m=1,0,-1$. In Fig. S2, we illustrate the form of the |$\ket{s}$ and |$\ket{t_{m}}$ for $J_{\perp}\gg J_{H}$ and $J_{\perp}\ll J_{H}$ limit. In Fig. S3(a,b), we plot the energy of the singlet and triplet states and their energy gap depending on $J_{\perp}/J_{H}$.
We now incorporate a finite in-plane exchange $J_{\parallel}$ by expressing the spin operators of the $d_{x^{2}-y^{2}}$ orbital in terms of the interlayer singlet and triplet basis. Within the restricted Hilbert space spanned by {|$\{\ket{s},\ket{t_{m}}\}$, the spin operators for the $d_{x^{2}-y^{2}}$ orbital ($\vec{S}^{d}$) is written in the form:
@@
[…]
@@
Here, $l=1,2$ denotes the layer, and $a=d,f$ labels the $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, respectively. The bosonic operators satisfy the canonical commutation relation $[b_{i,l,a},b^{\dagger}_{i^{\prime},l^{\prime},a^{\prime}}]=\delta_{i,i^{\prime}}\delta_{l,l^{\prime}}\delta_{a,a^{\prime}}$.
@@
where temperature dependence enters through the Bose–Einstein distribution $n_{\mathrm{BE}}(\Omega_{\vec{k},\eta,a}/T)=\langle\tilde{b}_{\vec{k},\eta,a}^{\dagger}\tilde{b}_{\vec{k},\eta,a}\rangle$. $\mathcal{N}$ is number of the sites. In Fig. S4, we plot the $(J_{H},J_{\perp})$ and $T$ dependence of $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$ by sustituting $S=1/2$ in Eq. (S33). For numerical integration over $\vec{k}$, we discretize the interval $k_{x},k_{y}\in[-\pi,\pi]$ using a step size of $2002\pi/$ 200.
Figure S4: Holstein-Primakoff spin-wave theory results with $S=1/2$ for $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$. Here, we fix $J=1$. (a,b) Zero-temperature $(J_{H},J_{\perp})$ dependence of the staggered moments $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$. The $d_{x^{2}-y^{2}}$ moment grows with increasing $J_{H}$ and decreasing $J_{\perp}$, while the $d_{z^{2}}$ moment is enhanced for small $J_{H}$ and small $J_{\perp}$. As $J_{H}\to 0$, $\langle N_{x^{2}-y^{2}}\rangle$ approaches 0.30, consistent with the known zero-temperature HP result for a square lattice with $S=1/2$. As $J_{H}\to 0$, $\langle N_{x^{2}-y^{2}}\rangle$ approaches 0.30, consistent with the HP result for a square lattice with $S=1/2$. In the $J_{H}\to\infty$ and small $J_{\perp}$ limit, $\langle N_{x^{2}-y^{2}}\rangle+\langle N_{z^{2}}\rangle\to 0.80$, matching the $S=1$ result. (c,d) Temperature dependence of $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$ at fixed $J_{\perp}=5$.
@@
The physical meaning of $N$, $dN$ is the orbital summed AFM order, and their differences.
@@
$E_{GS}(M)=\frac{1}{4}\sum_{n=1}^{8}\sum_{k_{z}=0,\pi}\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}E_{n}(\vec{k})\Theta[-E_{n}(\vec{k})]+a_{+}{N^{2}}$ (S51)

### v2

@@ Abstract
Introduction — Soon after the discovery of high temperature superconductivity (HTS) in a family of bilayer nickelates – first in pressurized bulk crystals [1, 2, 3, 4], and later in compressively strained films [5, 6, 7, 8] – intense effort has been devoted to understanding the emergent properties of these systems. Experiments have begun to shed some light on structural properties [9, 10], electronic band structure [11, 12, 13, 14] and magnetic orders proximate to superconductivity [15, 16, 17, 18, 19, 20, 21]. On the theoretical front, an impressive collection of works have already provided detailed proposals for the electronic structure, correlations and their implications for superconductivity [22, 23, 24, 25, 26, 11, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
In the wake of a new experimental discovery, there is the temptation towards classification, i.e. to ascertain whether the new phenomena belongs to a category of its own, or represents an interesting variation on a known theme. While HTS in the thin film bilayer nickelate is reminiscent of the overdoped cuprates [57], there remain obvious distinctions, the most prominent one being the discrepancy in valence count. For example, in the La-based cuprate parent compound La2CuO4, the low energy behavior is determined by the copper $d^{9}$ configuration, in which electronic excitations stem from the $d_{x^{2}-y^{2}}$ orbital and oxygen $p$ orbitals [58, 59, 60]. By contrast, the $d^{7.5}$ valence in the bilayer nickelate La3Ni2O7 necessitates the importance of multiple $d$-orbitals and Hund’s coupling. To determine whether such differences in local quantum chemistry lead to new emergent behavior, it is useful to take a step back and ask questions that are more qualitative in nature.
@@
Bilayer two-orbital Hubbard model — The bilayer nickelate consists of a $d^{7.5}$ configuration of Nickel electrons in a tetrahedral crystalline environment. Consequently, the low energy electronic excitations come from partially filled, non-degenerate Nickel $d_{3z^{2}-r^{2}}$ (denoted as Z``Z") and $d_{x^{2}-y^{2}}$ (denoted as X``X") orbitals. Thus, we begin with a two-orbital Hubbard model on a bilayer square lattice (see Fig. 1(a)), incorporating the $X,Z$ orbitals of Ni ions. Owing to the shape and orientation of these orbitals, it follows that electrons in the $Z$-orbital tunnel predominantly between the Nickel planes of the bilayer, whereas electrons in the $X$-orbital primarily tunnel within these planes (see Fig 1(a)). The Hamiltonian that incorporates only these strongest orbital hybridization effects is
$H$ $=-t_{\parallel}\sum_{\langle i,j\rangle,\ell,\sigma}\left(d^{\dagger}_{i\ell\sigma}d_{j\ell\sigma}+\text{h.c.}\right)-t_{\perp}\sum_{i,\sigma}\left(f^{\dagger}_{i1\sigma}f_{i2\sigma}+\text{h.c.}\right)$
@@
Magnetism in the insulating $d^{8}$ limit — At $x=0$, both the $X$ and $Z$-orbitals are half-filled and the model in Eq. (1) reduces to an effective two orbital, spin-1/2 bilayer Heisenberg model (see Fig. 2):
@@
Figure 3: (a) Phase diagram from bond-operator mean-field theory (BOMFT). Blue, yellow region denotes interlayer singlet and Néel order, respectively. The black line indicates the second order phase transition boundary between these two phases. (b) $J_{H}$ dependence of orbital-summed Néel order, $N=N_{x^{2}-y^{2}}+N_{z^{2}}$ (purple line) from Holstein-Primakov theory at $J_{\perp}=5$. Orbital-resolved Néel orders, $N_{x^{2}-y^{2}}=(-1)^{i+l}\langle S^{d;z}_{i,l}\rangle$ and $N_{z^{2}}=(-1)^{i+l}\langle S^{f;z}_{i,l}\rangle$, are plotted in blue, red dashed line, respectively. The system transitions from low-spin (spin-1/2) AFM on the $X$-orbital at small $J_{H}$ to high-spin (spin-1) AFM from both orbitals at large $J_{H}$. Inset plots $\partial N/\partial J_{H}$ which highlights that $N_{x^{2}-y^{2}}$ decreases while $N_{z^{2}}$ increases with increasing $J_{H}$ at large $J_{H}$ regime.
@@
Within the BOMFT, one begins by assuming an interlayer singlet ground state and describes the system using a local singlet–triplet basis, which corresponds to the ground and first excited states in the $J_{\parallel}\rightarrow 0$ limit. The critical interlayer coupling $J_{\perp}^{c}$ is then determined by the point at which the triplet excitation gap closes, due to the finite in-plane coupling $J_{\parallel}$ effects. Using two variational parameters—the condensation of interlayer singlet $\langle s\rangle$ and the chemical potential $\mu$—we numerically solve the self-consistent Hamiltonian and determine the phase boundary and its analytic form [67]. Fig. 3(a) shows the resulting phase diagram from BOMFT. As expected, $J_{\perp}^{c}=\infty$ at $J_{H}=0$ and it decreases with increasing $J_{H}$ and approaches $J_{\perp}^{c}\rightarrow 6.10$ in the $J_{H}\rightarrow\infty$ limit, in closer agreement with the QMC result $J_{\perp}^{c}=7.15$ for the bilayer spin-1 Heisenberg model [63]. Therefore, the overall range of $J_{\perp}^{c}$ estimated by BOMFT is $6.10<J_{\perp,c}^{\mathrm{BOMFT}}<\infty$. In contrast, the critical value obtained from the Schwinger-boson mean-field theory (SBMFT) lies in the range $13.5<J_{\perp,c}^{\mathrm{SBMFT}}<\infty$, which is significantly larger than the BOMFT estimate. This discrepancy reflects the well-known tendency of mean-field approaches to overestimate the stability of ordered phases. Specifically, BOMFT (SBMFT) starts from the interlayer-singlet (Néel-ordered) limit, and therefore tends to underestimate (overestimate) $J_{\perp,c}$. Consequently, the true critical value should lie between the two mean-field estimates, $J_{\perp,c}^{\mathrm{BOMFT}}<J_{\perp,c}<J_{\perp,c}^{\mathrm{SBMFT}}$. Moreover, in the large-$J_{H}$ regime, the BOMFT results are expected to be closer to the actual value, as tabulated in Table. S1, and they more faithfully capture the physics of the $_{\parallel}JJ_{\perp}$$J_{H}$ model.
@@
Next, to understand the spin-wave spectrum deep inside the antiferromagnetic phases, we use the standard Holstein–Primakov (HP) spin-wave theory [64, 68]. We evaluate the orbital-resolved Néel order $N_{x^{2}-y^{2}}=(-1)^{i+l}\langle S^{d;z}_{i,l}\rangle$ and $N_{z^{2}}=(-1)^{i+l}\langle S^{f;z}_{i,l}\rangle$ and their sum, $N\equiv N_{x^{2}-y^{2}}+N_{z^{2}}$. In Fig. 3(b), we plot $J_{H}$ dependence of $N$, $N_{x^{2}-y^{2}}$ and $N_{z^{2}}$ at a representative value $J_{\perp}=5$, where the AFM phase is expected in all range of $J_{H}$. For small $J_{H}$, $N_{z^{2}}$ remains suppressed and only $N_{x^{2}-y^{2}}$ exhibits finite magnetic order. This corresponds to a low-spin (spin-1/2) antiferromagnetic state. As $J_{H}$ increases, Hund’s coupling more transmits $J_{\perp}$ to the $X$-orbital, suppressing $N_{x^{2}-y^{2}}$, while simultaneously enhancing the transmission of $J_{\parallel}$ to the $Z$-orbital, thereby increasing $N_{z^{2}}$ and $N$. In the large-$J_{H}$ limit, the two order parameters converge to the same value, indicating a high-spin (spin-1) antiferromagnetic state. Even in the $J_{H}\to\infty$ limit with a small $J_{\perp}$, $N\to 0.80$, matching the $S=1$ HP result. In SM, we present a full map of the order parameters across the $(J_{\perp},J_{H})$ space [67]. By varying $J_{\perp}$, we find that the energy scale of $J_{H}$ associated with the crossover from the low-spin to the high-spin state decreases as $J_{\perp}$ is reduced. Thus, we have used several complementary approaches to establish that magnetism in the $d^{8}$ insulating regime can take on either high-spin, low spin forms, or can exhibit a spin-gapped, non magnetic ground state.
@@
Figure 4: Hartree-Fock (HF) results for Néel order (NN) under hole doping $x$ of (a) low-spin and (b) high-spin configurations for values of $U/t_{\parallel}$. Here, $N=(-1)^{i+\ell}\langle S^{d;z}_{i,\ell}+S^{f;z}_{i,\ell}\rangle$ is an orbital-summed Néel order. We used the hopping parameters $t_{\parallel}=0.483$, $t_{\perp}=0.635$, and onsite energy difference $\Delta=0.367$ from DFT calculations [22]. In both cases, we observe first-order or weakly first-order transitions. The AFM order in the high-spin case is more robust than in the low-spin case, as indicated by a larger critical doping $x_{c}$
We present the HF results for an intermediate $U=4t_{\|}$ in Fig. 4. In the low spin case (Fig. 4(a)), when the two orbitals are completely decoupled ($J_{H}=0$), the HF result coincides with the single layer single orbital square lattice Hubbard model, where for $U=4t$, AFM order vanishes at $x_{c}\approx 0.207$ through a first order phase transition [69]. Such first order antiferromagnetic phase transitions at $T=0$ are typical within the HF approximation. With a larger $J_{H}=0.4U$, the low spin AFM order gets suppressed compared to the decoupled limit: the spin order survives to a smaller $x_{c}\approx 0.18$ and its magnitude also becomes smaller. By contrast, the high spin AFM can survive to a much larger hole doping, as can be seen in Fig. 4(b). Moreover, the nature of the high spin AFM phase transition as a function of $x$ is also different from the low spin counterpart. We found that at large enough $J_{H}$, the high spin AFM order parameter $N$ can exhibit a discontinuous jump to a smaller value, which is due to the fact that the ground state energy has two local minima as a function of $N$, and increasing $x$ shifts the global minimum from one of two local minima to the other (see SM for details). But after this it appears to continuously decrease as $x$ increases further. This behavior seems to be a combination between both a first order and a continuous phase transition. Ref. [51], which employed a renormalized mean-field (Gutzwiller) approach starting from the type-II t-J model asuming infinite-Hund’s-coupling limit, reported a critical doping of $x_{c}\sim 0.55$ for the disappearance of antiferromagnetism (G-AFM). This value is consistent with our result high spin results, despite the different model assumptions.
@@
Finally, we comment on some limitations of our Hartree–Fock (HF) calculation. In this work, we have considered only the simple Néel order, which is naturally expected to dominate near the $d^{8}$ parent compound. However, closer to the $d^{7.5}$ filling relevant to bilayer nickelates, the Néel order may compete with other spin-density-wave (SDW) configurations such as $(\pi/2,\pi/2)$ or $(\pi,0)$ orders [56], which should be included in a more comprehensive mean-field ansatz. Another open issue is possible phase separation, which has been extensively discussed in the single-layer ttJJ model [70]. Investigating phase separation and other competing orders in the bilayer system remains an interesting direction for future work.
Discussion — We have analyzed parent compounds of bilayer nickelate with $d^{8}$ electron configuration and have shown that there are several distinct types of magnetic behavior depending on the interplay between Hund’s coupling ($J_{H}$) and superexchange ($J_{\bot},J_{\|}$). We have identified a crossover from a low-spin AFM order at small $J_{H}$ to a high-spin AFM order at larger $J_{H}$. We have also identified a spin-gapped insulating state that occurs when $J_{\perp},J_{H}\gg J_{\parallel}$. An important remark concerns the parameter regime $J_{\perp},J_{H}\gg J_{\parallel}$ relevant to realistic materials. While some DMFT-based analyses have estimated $J_{H}/J_{\parallel}\approx 5$ and $J_{\perp}/J_{\parallel}\sim 2$ for bilayer nickelates La3Ni2O7 ($d^{7.5}$ configuration), the exchange couplings are expected to evolve with doping $x$, and their realistic magnitudes in the $d^{8}$ limit remain an open question.
The nature of magnetism in the parent compound will certainly have implications for superconductivity. We have shown here that the low energy physics of the doped low-spin state is adequately described by a single orbital bilayer Hubbard model when $J_{H}/J_{\perp}$ is vanishingly small. When $J_{H}/J_{\perp}$ is small, but finite, an exchange coupling between the two layers will be generated by virtual processes. In contrast, the low energy effective description of the doped high spin state is a bilayer-Kondo lattice with ferromagnetic Kondo coupling due to Hund’s coupling and a weak interlayer exchange coupling. If one considers the extreme limit where $J_{H}\rightarrow\infty$, the Kondo model can be further simplified to a bilayer type-II ttJJ model, as two of us derived in the previous paper [24]. We stress that all three different regimes can be reconciled by a double-Kondo model  [62].
Within Hartree-Fock mean-field theory, we find that upon hole doping, the high-spin antiferromagnetic state becomes more robust. From this observation, it remains unclear how such robust magnetism influences superconductivity. Two opposite scenarios can be made: in one case, the enhanced magnetic order competes more strongly with superconductivity and may suppress pairing; in the other, within a spin-fluctuation framework, stronger magnetism could instead provide a more effective pairing ‘glue’ than in the low-spin counterpart. Lastly, the spin gapped insulator, when doped exhibits superconducivity rather distinct than either the high- or low-spin states, with strong interalyer pairing as was shown recently [61, 62].
@@
Our findings are pertinent to recent experimental studies with compressively strained thin films of the bilayer nickelate. In particular, we encourage future experiments to explore the $d^{8}$ insulating limit and to probe possible magnetic transitions via strain engineering, which can be effectively described by varying $J_{\perp}/J_{\parallel}$ at fixed $J_{H}$. With compressive strain, we naturally expect that the in-plane superexchange will be enhanced whereas with tensile strain, the opposite would hold. Given the rich magnetic phase diagram revealed in our study, the experimental observation of distinct spin states in strain-tuned systems represents a promising arena for future exploration.
@@ 5 Departments of Applied Physics, Stanford University, Stanford, CA 94305, USA
$H_{MF}$ $=$ $\frac{J_{\parallel}}{4}\sum_{\langle i,j\rangle,l,\sigma}\Delta_{\parallel}b_{i,l,d,\sigma}^{\dagger}b_{j,l,d,\sigma}^{\dagger}+\text{h.c.}+J_{\perp}\sum_{i,\sigma}\Delta_{\perp}b_{i,1,f,\sigma}^{\dagger}b_{i,2,f,\sigma}^{\dagger}+\text{h.c.}$ (S7)
@@
In our numerical implementation, we approximate the momentum summation $\sum_{k}=\int d\gamma,D(\gamma)$ by treating the density of states $D(\gamma)$ as a constant value of $1/2$, following the standard approximation used in prior literature [72]. For numerical integration over $\gamma$, we discretize the interval $\gamma\in[-1,1]$ using a step size of $1/2000$.
@@
[…]
@@
$+C_{st}^{2}(r)\,\overline{s}^{2}\sum_{\langle i,j\rangle}\left[(t_{i,0}+t^{\dagger}_{i,0})(t_{j,0}+t^{\dagger}_{j,0})+(t_{i,-1}-t^{\dagger}_{i,1})(t_{j,-1}-t^{\dagger}_{j,1})+\text{h.c.}\right],$
@@
Here, $l=1,2$ denotes the layer, and $a=d,f$ labels the $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, respectively. The bosonic operators satisfy the canonical commutation relation $b_{i,l,a},b^{\dagger}_{i^{\prime},l^{\prime},a^{\prime}}]=\delta_{i,i^{\prime}}\delta_{l,l^{\prime}}\delta_{a,a^{\prime}}$.
@@
where temperature dependence enters through the Bose–Einstein distribution $n_{\mathrm{BE}}(\Omega_{\vec{k},\eta,a}/T)=\langle\tilde{b}_{\vec{k},\eta,a}^{\dagger}\tilde{b}_{\vec{k},\eta,a}\rangle$. $\mathcal{N}$ is number of the sites. In Fig. S4, we plot the $(J_{H},J_{\perp})$ and $T$ dependence of $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$ by sustituting $S=1/2$ in Eq. (S33). For numerical integration over $\vec{k}$, we discretize the interval $k_{x},k_{y}\in[-\pi,\pi]$ using a step size of $2\pi/200$.
Figure S4: Holstein-Primakoff spin-wave theory results with $S=1/2$ for $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$. Here, we fix $J=1$. (a,b) Zero-temperature $(J_{H},J_{\perp})$ dependence of the staggered moments $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$. The $d_{x^{2}-y^{2}}$ moment grows with increasing $J_{H}$ and decreasing $J_{\perp}$, while the $d_{z^{2}}$ moment is enhanced for small $J_{H}$ and small $J_{\perp}$. As $J_{H}\to 0$, $\langle N_{x^{2}-y^{2}}\rangle$ approaches 0.30, consistent with the known zero-temperature HP result for a square lattice with $S=1/2$. As $J_{H}\to 0$, $\langle N_{x^{2}-y^{2}}\rangle$ approaches 0.30, consistent with the HP result for a square lattice with $S=1/2$. In the $J_{H}\to\infty$ and small $J_{\perp}$ limit, $\langle N_{x^{2}-y^{2}}\rangle+\langle N_{z^{2}}\rangle\to 0.80$, matching the $S=1$ result. (c,d) Temperature dependence of $\langle N_{x^{2}-y^{2}}\rangle$ and $\langle N_{z^{2}}\rangle$ at fixed $J_{\perp}=5$. (e) Order parameter at different $J_{\perp}$. This demonstrates that the energy scale of $J_{H}$ associated with the crossover from the low-spin to the high-spin state decreases as $J_{\perp}$ is reduced.
@@
The physical meaning of $N$, dNdN is the orbital summed AFM order, and their differences.
@@
$E_{GS}(N)=\frac{1}{4}\sum_{n=1}^{8}\sum_{k_{z}=0,\pi}\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}E_{n}(\vec{k})\Theta[-E_{n}(\vec{k})]+a_{+}{N^{2}}$ (S51)
@@
An example of the ground state energy is shown in Fig.S6. Reading the global minimum from these curves determines Fig.4(b) in the main text. Here we only show three curves near the vicinity of the discontinuity shown in Fig.4(b) as the sudden jump of $N$. From Fig.S6 we clearly see that the such a sudden jump is intimately related to the presence of two local minima in the energy profile. As $x$ increase, the global minimum shifts from one of the two local minima to the other in a first order transition manner.
Figure S6: Ground state energy as a function of $N$ for the high-spin case. Here we only present three curves near the critical doping where the discontinuous jump occurs. It is apparent that the discontinuity in $N$, i.e. the global minimum, originates from the fact that there are two local mimina in the energy profile.